OpenStudy (jpsmarinho):

How to integrate this: Integrate(x^2*e^x) with dx , starting at 0 to 1?

OpenStudy (anonymous):

Integrate by parts twice.

OpenStudy (anonymous):

put \[ u =x^2; \quad dv=e^x dx\\ du= 2 x; v= e^x\\ \int x^2 e^x dx = x^2 e^x - \int 2 x e^x dx \]

OpenStudy (anonymous):

For the last integral put \[ u=2x ;\quad dv=e^x\\ du= 2 dx ; \quad v=e^x \int 2 x e^x dx = 2 xe^x - \int 2 e^x dx=\\ 2 xe^x - 2 e^x \]

OpenStudy (anonymous):

Combining and simplifying, you get \[ \int x^2 e^x \, dx=e^x \left(x^2-2 x+2\right) \]

OpenStudy (anonymous):

\[ \int_0^1 x^2 e^x \, dx=\left[e^x \left(x^2-2 x+2\right)\right]_0^1=-2 +e \]

OpenStudy (jpsmarinho):

Thanks!